Charges of +2.0, +3.0, and -8.0C are placed at the vertices of an equilateral triangle of side 10 cm. Calculate the magnitude of the force acting on the -8.0C charge due to the other two charges.

1
Expert's answer
2022-01-26T17:45:56-0500
"F_x=\dfrac {q\cos30°}{4\pi\epsilon_0r^2}(q_1+q_2),\\\space\\
F_y=\dfrac {q\sin30°}{4\pi\epsilon_0r^2}(q_1+q_2),\\\space\\
F_\text{net}=\sqrt{F_x^2+F_y^2}=\dfrac {q(q_1+q_2)\sqrt2}{4\pi\epsilon_0r^2}=5.08·10^{13}\text{ N}."